# Extracting Probabilistic Knowledge from Large Language Models for Bayesian Network Parameterization

### v1

@@ 7 Conclusion and Future Work
In conclusion, this study introduces a novel pipeline for parameterizing BNs using LLMs and establishes the first comprehensive framework for evaluating the probabilistic knowledge of LLMs with real-world probabilistic BNs. We systematically assess various BNs and demonstrate that modern LLMs can reliably produce conditional probabilities, outperforming baselines and purely data-driven approaches, especially in a low-data regime. Importantly, our evaluation framework highlights the probabilistic estimation capabilities of LLMs, offering insights into their internal knowledge, demonstrating their value as expert-based priors, and facilitating automated BN construction.
Future work will focus on advancing toward a fully automated framework for Bayesian Network construction. In this regard, the key challenge lies in automating the structure learning component. Although preliminary efforts have been made in this area, there is significant potential to create an end-to-end pipeline using LLMs that generates a BN structure, parameterizes it, and systematically evaluates its performance.
Limitations
Our analysis demonstrates clear advantages in employing LLMs for Bayesian network parameter estimation. While KL divergence is effective for comparative evaluation within our context, it primarily indicates relative improvements rather than quantifying specific practical benefits for individual downstream tasks, which often require tailored metrics, such as assessing the accuracy of query probabilities within a defined precision (e.g., a 5% threshold). Incorporating domain expert judgments to evaluate the BNs and design specific metrics for each BN would benefit our work, but the associated costs and resource requirements substantially limit this approach. Expert evaluations would realistically be feasible only for a select subset of Bayesian networks rather than all 80 networks analyzed, limiting the comprehensiveness and comparative scope of the evaluation currently achievable through automated methods.
Additionally, our study utilizes only state-of-the-art models because smaller, more economical models performed poorly in the task of probabilistic knowledge estimation. Although our approach avoids fine-tuning and thus has lower costs than fine-tuning-based methods, the inference costs remain relatively high among inference-only approaches, potentially posing challenges in particularly cost-sensitive scenarios.
@@ Appendix A Dataset Pre-processing
A.1 bnRep Dataset Overview
The bnRep dataset (Leonelli, 2025) is an open-source collection designed to facilitate research, teaching, and practical applications related to Bayesian Networks (BNs), addressing the significant shortage of comprehensive BN repositories. Implemented as an R package, bnRep includes over 200 well-documented Bayesian Networks sourced from more than 150 academic publications, mainly recent studies published from 2020 onwards. Each BN entry has accompanying characteristics extracted from the original publications. These characteristics are described below:
@@
A.2 BN Selection and Filtering
To utilize the bnRep dataset, we first converted the Bayesian networks from the R package into Python-compatible format using the pgmpy library (Ankan and Panda, 2015). Next, we filtered the networks, selecting only those containing discrete CPT values, as these comprise the majority (90%) of the BNs. Then, to improve practicality and computational efficiency, we excluded networks with more than 50 nodes, noting that only a few exceed this threshold. The most substantial reduction, however, came from removing BNs with incomplete CPT information. After applying all these criteria, we arrived at a final dataset of 80 BNs, which remains sufficiently large for our evaluation purposes.
A.3 Nodes’ Descriptions and States
As outlined in Section 4, querying the LLMs for each node state requires clearly defined descriptions for the nodes and their respective states. To acquire this information for each Bayesian Network (BN), we first retrieved the PDFs of the referenced documents that detail each BN. These references were initially provided by the bnRep dataset.
A Python script was developed to automate the download of relevant PDFs directly from Google Scholar using bibliographic information. Despite automation, several documents could not be downloaded automatically due to access restrictions or other limitations. These remaining documents were obtained manually, with some requiring specific permissions. Subsequently, we performed a manual verification to ensure the downloaded documents corresponded to their respective references.
We then implemented another Python script leveraging GPT-4o, which, given a PDF and the extracted nodes and states of the Bayesian Network (obtained through the pgmpy library), automatically generated a Python dictionary describing each node and its associated states. Although the majority of generated dictionaries were accurate and required minimal modifications, a few necessitated manual adjustments to ensure correct formatting and accuracy. An illustrative example of the prompt and the generated dictionary for the Bayesian Network “algalactivity2” is provided in Figure 4.
@@ A.3 Nodes’ Descriptions and States
Appendix B LLM’s Hyper-Parameters and Experimental Setup
In our experiments, we primarily evaluated three state-of-the-art Large Language Models: GPT-4o and its mini variant (OpenAI et al., 2024) versions “gpt-4o-2024-11-20” and “gpt-4o-mini-2024-07-18”, Claude Sonnet 3.5 (Anthropic, 2024) version “Claude 3.5 Sonnet 2024-10-22”, and Gemini-Pro 1.5 (Team et al., 2024) version “gemini-1.5-pro-002”. All models were interfaced using the LangChain framework (Chase, 2022), ensuring consistent integration and interaction.
A fixed generation “temperature” of $0.1$ was utilized to maintain minimal stochasticity in outputs. We initially explored the impact of sampling by performing up to five repeated samples per inference. However, we found that multiple samples did not meaningfully affect aggregate outcomes across the evaluated set of 80 Bayesian Networks, likely due to the low temperature setting. Therefore, subsequent experiments were conducted using a single sample per inference.
Token output lengths were not constrained, allowing the models to elaborate their reasoning freely. In instances where models produced responses that deviated from the required format or where output text generation was interrupted midway, additional prompts were provided until valid responses were obtained. The prompt structure used for eliciting responses is illustrated at the top of Figure 5. To get the token probabilities from our LLMs, we use the middle prompt template in Figure 5. OpenAI’s Chat API allows for extracting the top 20 most probable tokens. To get the full distribution of a node’s states at once using the Full Dist method, we use the prompt template at the bottom of Figure 5.
Figure 5: Prompt templates used for eliciting probabilistic responses from LLMs. The top panel shows the main prompt template that asks the LLM to provide a probabilistic answer along with explicit reasoning. The middle panel shows the modified prompt template designed specifically to extract token-level probabilities by instructing the LLM to generate only the most probable state without additional context or explanation. The bottom panel shows the prompt template used for the Full Dist method, which requests the model to output the entire probability distribution over a node’s possible states in a single response.
Appendix C KL Divergence
C.1 Kullback Leibler (KL) Divergence Overview
Kullback Leibler (KL) divergence measures how one probability distribution $p$ diverges from a reference distribution $q$. For a discrete random variable $X$,
@@ C.1 Kullback Leibler (KL) Divergence Overview
For continuous variables the sum is replaced by an integral. KL divergence satisfies $D_{\mathrm{KL}}(p\|q)\geq 0$ and equals zero iff $p=q$. It is asymmetric ($D_{\mathrm{KL}}(p\|q)\neq D_{\mathrm{KL}}(q\|p)$ in general). Intuitively, it quantifies the expected extra amount of information (in nats or bits) required to encode samples from $p$ using a code that is optimal for $q$.
C.2 BN KL Divergence Calculation
@@
Let $A,B,C\in\{0,1\}$ within the network $A\!\to\!B\!\to\!C$. We compute $D_{\mathrm{KL}}(p\|q)$ as follows: 1. Substitute the factorizations: $D_{\mathrm{KL}}(p\|q)=\sum_{a,b,c}p(a,b,c)\,\log\frac{p(a,b,c)}{q(a,b,c)}.=\sum_{a,b,c}p(a,b,c)\,\log\frac{p(a)\,p(b\mid a)\,p(c\mid b)}{q(a)\,q(b\mid a)\,q(c\mid b)}.$ 2. Separate logs: $=\sum_{a,b,c}p(a,b,c)\Bigl{[}\log\frac{p(a)}{q(a)}+\log\frac{p(b\mid a)}{q(b\mid a)}+\log\frac{p(c\mid b)}{q(c\mid b)}\Bigr{]}.$ 3. Split the sum: $=\underbrace{\sum_{a,b,c}p(a,b,c)\,\log\frac{p(a)}{q(a)}}_{A}+\underbrace{\sum_{a,b,c}p(a,b,c)\,\log\frac{p(b\mid a)}{q(b\mid a)}}_{B}+\underbrace{\sum_{a,b,c}p(a,b,c)\,\log\frac{p(c\mid b)}{q(c\mid b)}}_{C}.$ 4. Marginalize: $A=\sum_{a}p(a)\,\log\frac{p(a)}{q(a)},\quad B=\sum_{a}p(a)\sum_{b}p(b\mid a)\,\log\frac{p(b\mid a)}{q(b\mid a)},\quad C=\sum_{b}p(b)\sum_{c}p(c\mid b)\,\log\frac{p(c\mid b)}{q(c\mid b)}.$ 5. Recognize KL pieces and combine: $D_{\mathrm{KL}}(p\|q)=D_{\mathrm{KL}}\!\bigl{(}p(A)\|q(A)\bigr{)}+\sum_{a}p(a)\,D_{\mathrm{KL}}\!\bigl{(}p(B\mid a)\|q(B\mid a)\bigr{)}+\sum_{b}p(b)\,D_{\mathrm{KL}}\!\bigl{(}p(C\mid b)\|q(C\mid b)\bigr{)}.$
@@
Appendix D Additional Diagrams
The results of the priors as pseudocounts method are shown in Figure 7. The trend and conclusions drawn from this figure are the same as the linear pooling method shown in Figure 3. Figure 8 shows the results of Figure 2 but with CPT KL divergence metric. The results of EDP with data sampled for each row of the CPT by # times the number of states of the node are shown in Figures 9 and 10.
Figure 7: Distribution of BN KL Divergence across 80 Bayesian Networks using GPT-4o (EDP) and Uniform predictions as priors combined with empirical data via the priors as pseudocounts method.
@@ Appendix D Additional Diagrams
Figure 8: Boxplot showing the distribution of BN KL Divergence values across 80 unique Bayesian Networks for various models and baselines. Models are sorted by their median KL Divergence, with lower values indicating better alignment with ground truth CPTs.
Figure 9: Distribution of BN KL Divergence across 80 Bayesian Networks using GPT-4o (EDP) and Uniform predictions as priors combined with empirical data via the priors as pseudocounts method.
Figure 10: Distribution of BN KL Divergence across 80 Bayesian Networks using GPT-4o (EDP) and Uniform predictions as priors combined with empirical data via the linear pooling method.
Appendix E Differnt number parnets and states
Figures 11, 12, 13, 14 and 15 show the CPT KL divergence for nodes with 2, 3, 4, 5 and 6 states respectively. Figures 16, 17, 18, 19, 20, 21, 22 and 23 show the CPT KL divergence for nodes with 0, 1, 2, 3, 4, 5, 6 and 7 parents respectively.

### v2

@@ 7 Conclusion and Future Work
In this work, we demonstrate that modern LLMs can effectively produce conditional probabilities to be used as expert-driven priors for Bayesian Networks parameterization. EDP proved superior to the approach of using Uniform priors and purely data-driven approaches, especially in a low-data regime. Furthermore, we show that employing EDP to estimate parameters of the BNs improved the accuracy in downstream tasks. In conclusion, this study introduces a novel pipeline for parameterizing BNs using LLMs. Using LLMs as a resource compensates for the lack of data and reduces the need for costly domain experts. We also establish the first comprehensive framework for evaluating the probabilistic knowledge of LLMs with real-world probabilistic BNs.
Future work will focus on advancing toward a fully automated framework for BN construction. In this regard, the key challenge lies in automating the structure learning component. Although preliminary efforts have been made in this area, there is significant potential to create an end-to-end pipeline using LLMs that generates a BN structure, parameterizes it, and systematically evaluates its performance.
@@ Appendix A Dataset Pre-processing
bnRep Dataset Overview
The bnRep dataset (Leonelli 2025) is an open-source collection designed to facilitate research, teaching, and practical applications related to Bayesian Networks (BNs), addressing the significant shortage of comprehensive BN repositories. Implemented as an R package, bnRep includes over 200 well-documented Bayesian Networks sourced from more than 150 academic publications, mainly recent studies published from 2020 onwards. Each BN entry has accompanying characteristics extracted from the original publications. These characteristics are described below:
@@
BN Selection and Filtering
To utilize the bnRep dataset, we first converted the Bayesian Networks from the R package into Python-compatible format using the pgmpy library (Ankan and Panda 2015). Next, we filtered the networks, selecting only those containing discrete CPT values, as these comprise the majority of the BNs. Then, to improve practicality and computational efficiency, we excluded networks with more than 50 nodes due to practical and cost considerations, noting that only a few exceed this threshold. The most substantial reduction, however, came from removing BNs with incomplete CPT information. After applying all these criteria, we arrived at a final dataset of 80 BNs, which remains sufficiently large for our evaluation purposes.
Nodes’ Descriptions and States
Querying the LLMs for each node requires clearly defined descriptions for the nodes and their respective states. To acquire this information for each BN, we first retrieved the PDFs of the referenced documents that detail each BN. These references were initially provided by the bnRep dataset. A Python script was developed to automate the download of relevant PDFs directly from Google Scholar using bibliographic information. Despite automation, several documents could not be downloaded automatically due to access restrictions or other limitations. As a result, these remaining documents were obtained manually. Subsequently, we performed a manual verification to ensure the documents corresponded to their respective references.
We then implemented another Python script leveraging GPT-4o, which, given a PDF and the extracted nodes and states of the Bayesian Network (obtained through the pgmpy library), automatically generated a Python dictionary describing each node and its associated states. The majority of generated dictionaries were accurate and required minimal modifications; a few necessitated manual adjustments to ensure correct formatting and accuracy. An illustrative example of the prompt and the generated dictionary for the Bayesian Network “algalactivity2” is provided in Figure 4.
@@
Appendix B KL Divergence
Kullback Leibler (KL) Divergence Overview
Kullback-Leibler (KL) divergence (Kullback and Leibler 1951) measures how one probability distribution $p$ diverges from a reference distribution $q$. For a discrete random variable $X$,
@@ Appendix B KL Divergence
For continuous variables, the sum is replaced by an integral. KL divergence satisfies $D_{\mathrm{KL}}(p\|q)\geq 0$ and equals zero iff $p=q$. It is asymmetric ($D_{\mathrm{KL}}(p\|q)\neq D_{\mathrm{KL}}(q\|p)$ in general). Intuitively, it quantifies the expected extra amount of information (in nats or bits) required to encode samples from $p$ using a code that is optimal for $q$. In our implementation, all computations use base-2 logarithms (bits). To avoid undefined terms when $q(x)=0$ but $p(x)>0$, we apply elementwise $\varepsilon$-smoothing to both $p$ and $q$ with $\varepsilon=10^{-8}$:
$\tilde{r}(x)\;=\;\frac{r(x)+\varepsilon}{\sum_{x^{\prime}\in\mathcal{X}}\!\bigl{(}r(x^{\prime})+\varepsilon\bigr{)}}\quad\text{for }r\in\{p,q\},$
[…]
BN KL Divergence Calculation
@@
Let $A,B,C\in\{0,1\}$ within the network $A\!\to\!B\!\to\!C$. We compute $D_{\mathrm{KL}}(p\|q)$ as follows: 1. Substitute the factorizations: $D_{\mathrm{KL}}(p\|q)=\sum_{a,b,c}p(a,b,c)\,\log\frac{p(a,b,c)}{q(a,b,c)}.=\sum_{a,b,c}p(a,b,c)\,\log\frac{p(a)\,p(b\mid a)\,p(c\mid b)}{q(a)\,q(b\mid a)\,q(c\mid b)}.$ 2. Separate logs: $\sum_{a,b,c}p(a,b,c)\Bigl{[}\log\frac{p(a)}{q(a)}+\log\frac{p(b\mid a)}{q(b\mid a)}+\log\frac{p(c\mid b)}{q(c\mid b)}\Bigr{]}.$ 3. Split the sum: $\underbrace{\sum_{a,b,c}p(a,b,c)\,\log\frac{p(a)}{q(a)}}_{A}+\underbrace{\sum_{a,b,c}p(a,b,c)\,\log\frac{p(b\mid a)}{q(b\mid a)}}_{B}+\underbrace{\sum_{a,b,c}p(a,b,c)\,\log\frac{p(c\mid b)}{q(c\mid b)}}_{C}.$ 4. Marginalize: $A=\sum_{a}p(a)\,\log\frac{p(a)}{q(a)},\quad B=\sum_{a}p(a)\sum_{b}p(b\mid a)\,\log\frac{p(b\mid a)}{q(b\mid a)},\quad C=\sum_{b}p(b)\sum_{c}p(c\mid b)\,\log\frac{p(c\mid b)}{q(c\mid b)}.$ 5. Recognize KL pieces and combine: $D_{\mathrm{KL}}(p\|q)=D_{\mathrm{KL}}\!\bigl{(}p(A)\|q(A)\bigr{)}+\sum_{a}p(a)\,D_{\mathrm{KL}}\!\bigl{(}p(B\mid a)\|q(B\mid a)\bigr{)}+\sum_{b}p(b)\,D_{\mathrm{KL}}\!\bigl{(}p(C\mid b)\|q(C\mid b)\bigr{)}.$
@@
CPT KL Divergence Calculation
In some of our experiments, we report a CPT KL divergence defined over local CPT rows instead of the entire BN. For each node $X_{i}$ with parent set $pa_{i}$ and parent configuration set $\mathcal{P}_{i}$, define the local divergence
$d_{i,u}\;=\;D_{\mathrm{KL}}\!\big{(}p(X_{i}\mid pa_{i}{=}u)\,\|\,q(X_{i}\mid pa_{i}{=}u)\big{)},\quad u\in\mathcal{P}_{i}.$
Our CPT KL divergence is the unweighted mean over all CPT rows in the BN:
$\mathrm{\text{CPT KL divergence}}\;=\;\frac{1}{\sum_{i=1}^{n}|\mathcal{P}_{i}|}\sum_{i=1}^{n}\;\sum_{u\in\mathcal{P}_{i}}d_{i,u}.$
This metric treats every CPT row equally and is therefore suitable for subgroup analyses, e.g., by number of states or by number of parents.
Appendix C Setup and Hyper-Parameters
Large Language Models’ Versions
In our experiments, we evaluated multiple state-of-the-art Large Language Models: GPT-4o and its mini variant (OpenAI et al. 2024) versions “gpt-4o-2024-11-20” and “gpt-4o-mini-2024-07-18”, Claude 3.5 Sonnet (Anthropic 2024) version “Claude 3.5 Sonnet 2024-10-22”, Gemini-Pro 1.5 (Gemini Team et al. 2024) version “gemini-1.5-pro-002” and DeepSeek-V3 (DeepSeek-AI et al. 2025b), version “DeepSeek-V3-0324”. To test the reasoning models we used o3 and its mini variant (OpenAI 2025), versions “o3-2025-04-16” and “o3-mini-2025-01-31”, and DeepSeek-R1 (DeepSeek-AI et al. 2025a), version “DeepSeek-R1-0528”. All models were interfaced using the LangChain framework (Chase 2022), ensuring consistent interaction.
LLMs’ Hyper-Parameters
A “temperature” of $0.1$ was utilized to maintain minimal stochasticity in outputs. For reasoning models, the “reasoning effort” was set to medium. We initially explored the impact of sampling by performing up to five repeated samples per inference. However, we found that multiple samples did not meaningfully affect aggregate outcomes across the evaluated set of 80 Bayesian Networks, likely due to the low temperature setting. Consequently, given the number of LLMs and the dataset size, all subsequent experiments used a single sample to control cost. Output lengths were not constrained, allowing the models to elaborate their reasoning freely. In instances where models produced responses that deviated from the required format or where output text generation was interrupted midway, additional prompts were provided until valid responses were obtained.
EDP’s Hyper-Parameters
In our EDP formula, we have to select the hyper-parameter, $\alpha$, in the following formula:
$p_{i}=\frac{\alpha q_{i}+c_{i}}{\alpha+\sum_{j=1}^{m}c_{j}},\qquad i=1,\dots,m.$
where $q_{i}$ is the prior probability (from the LLM) and $c_{i}$ is the observed count for state $i$. Ideally, $\alpha$ is selected by testing various $\alpha$ values on a dev dataset of a downstream task. In our EDP experiments, where we are not testing a downstream task, we use a heuristic to select the $\alpha$. It is intuitive that as the number of data samples increases, the importance of priors decreases. As a result, we set $\alpha$ to be proportional to the inverse of the number of data samples.
For classification experiments using EDP, for a dataset with $N$ data samples, we set $\alpha=0.5\times N$ in the full-data regime. In the low-data regime, we select the optimal $\alpha$ from $0.5\times N,1.0\times N,2.0\times N$} using the remaining training data as a development set.
Figure 6: Prompt templates used for eliciting probabilistic responses from LLMs. The prompting structures for SepState, FullDist, and Extracting token probabilities are shown in the top, middle, and bottom panels, respectively.
LLMs’ Prompting Structure
Figure 6 shows the exact prompt we use to query the LLMs. All templates begin with a compact list of node names and natural-language descriptions (including each node’s state space) and then pose a probabilistic query conditioned on a specific parent configuration. For SepState, as shown in the top panel, for each state of the node, the model is asked a single question and instructed to provide a probability. We parse the trailing numeric value and repeat this for all states. For FullDist, as shown in the middle panel, the query requests the entire distribution in one response and explicitly supplies the order of states. We parse the tuple in the given order. For token probability, as shown in the bottom panel, the LLM is asked for only the most probable state name without any extra text. From the returned token scores for each candidate string, we compute the probabilities of the states of the nodes. If a node’s state is not among the 20 returned candidate strings, we assign it a probability of 0. This occurs in only 12% of cases for GPT-4o and GPT-4o-mini. In our experiments, the probabilities of the returned strings typically sum to approximately 99%, justifying our assumption of a 0% probability for the absent states.
GPT-4o
Hill Climbing Naïve Bayes
Dataset Full Data 20 Samples 10 Samples Full Data 20 Samples 10 Samples
SepState MLE EDP MLE EDP MLE EDP SepState MLE EDP MLE EDP MLE EDP
[…]
DeepSeek-V3
Hill Climbing Naïve Bayes
Dataset Full Data 20 Samples 10 Samples Full Data 20 Samples 10 Samples
FullDist MLE EDP MLE EDP MLE EDP FullDist MLE EDP MLE EDP MLE EDP
[…]
Table 2: Macro-F1 classification scores on nine datasets for BN classifiers. Columns include three methods, SepState or FullDist using GPT-4o or DeepSeek-V3, MLE, and EDP, under two graph structures, HillClimbing and Naïve Bayes. Full Data uses the entire training split, whereas 20 and 10 Samples simulate low-resource regimes. Within each data regime, the higher of MLE vs. EDP is bolded. * datasets HV84, PhDA and Traject refer to HouseVote84, PhDArticles and Trajectories, respectively.
Appendix D Additional Experiments
EDP’s Impact on Downstream Tasks
Table 2 expands the classification study described in the main text by reporting additional results for DeepSeek-V3 and showing the variance in the results obtained in multiple runs. In the low-resource settings, the training cases are sampled from the training split, and results are averaged over five random runs. The $\pm$ values in the 20 and 10 sample columns denote the empirical standard deviation across the five runs. For the LLM-only reference column, we use SepState with GPT-4o and FullDist with DeepSeek-V3. The corresponding EDP columns use the same extraction scheme as their LLM prior. These detailed results, which include DeepSeek-V3 in addition to GPT-4o, support the claims made in the main text: across both structures and most datasets, EDP matches or improves upon MLE, with the largest gains in the low-resource regimes.
EDP with Other LLMs and Sampling Methods
In this paper, we demonstrated that EDP improves probability estimates as measured by KL divergence, using GPT-4o. However, this trend is consistent across other LLMs as well. Figures 7, 8, and 9 illustrate EDP’s results for DeepSeek-V3, Gemini-pro 1.5, and Claude 3.5 Sonnet, respectively.
Figure 7: Boxplot of distribution of BN KL divergence over 80 networks, contrasting models with DeepSeek-V3 priors (EDP‑#), uniform priors (Uniform‑#), and data‑only estimates (MLE‑#). The numeral after the “‑” denotes the sample size.
Figure 8: Boxplot of distribution of BN KL divergence over 80 networks, contrasting models with Gemini-pro 1.5 priors (EDP‑#), uniform priors (Uniform‑#), and data‑only estimates (MLE‑#). The numeral after the “‑” denotes the sample size.
Figure 9: Boxplot of distribution of BN KL divergence over 80 networks, contrasting models with Claude 3.5 Sonnet priors (EDP‑#), uniform priors (Uniform‑#), and data‑only estimates (MLE‑#). The numeral after the “‑” denotes the sample size.
In our main experiments, we performed forward sampling on the entire BN. However, EDP consistently improves results under alternative sampling methods. Figure 10 shows EDP’s performance when sampling # data points per CPT row. Similarly, Figure 11 presents the results when using a related sampling strategy, where each CPT row is sampled $\times n$ times, with $n$ representing the number of states of the node. These results prove that EDP’s improvements are independent of the sampling method.
Figure 10: Boxplot of distribution of BN KL divergence over 80 networks, contrasting models with GPT‑4o priors (EDP‑#), uniform priors (Uniform‑#), and data‑only estimates (MLE‑#). The numeral after the “‑” denotes the sample size. In this figure, the data for each row of the CPT is sampled by # instances.
Figure 11: Boxplot of distribution of BN KL divergence over 80 networks, contrasting models with GPT‑4o priors (EDP‑#), uniform priors (Uniform‑#), and data‑only estimates (MLE‑#). The numeral after the “‑” denotes the sample size. In this figure, the data for each row of the CPT is sampled by $\times n$ times, with $n$ representing the number of states of the node.
SepState and FullDist with CPT KL Divergence
Figure 12 reports CPT KL for the methods discussed in the main paper. Consistent with the BN KL divergence, we observe that the relative ordering of methods is unchanged, meaning that the SepState performs the best among all baselines. Here, however, FullDist performs worse than the Uniform baseline in most LLMs.
@@ Appendix D Additional Experiments
Figure 12: Boxplot showing the distribution of CPT KL divergence values across 80 unique BNs for various models, sorted by their median KL divergence. Lower values indicate better alignment with ground truth CPTs.
Appendix E Varying Number of Parents and States
Figures 13 through 17 illustrate the CPT KL divergence for nodes with 2 to 6 states, respectively. Figures 18 through 25 depict the CPT KL divergence for nodes with 0 to 7 parents, respectively. These figures help to show the points made in the discussion section about the capability of LLMs in handling nodes with various state sizes and parents.